What is the magnitude of the electric field 38.1 cm directly above an isolated 3.50×10−5 C charge?
Answer #1

Charge, q = 3.5 * 10^-5 C
r = 38.1 cm = 0.381 m

E = k*q1/r^2
E = (8.9*10^9*3.5 * 10^-5)/(0.381)^2
E = 2.15 * 10^6 N/C
Magnitude of the electric field, E = 2.15 * 10^6 N/C